# DATABench: Evaluating Dataset Auditing in Deep Learning from an Adversarial Perspective

### v1

@@ IV-C Forgery Attacks
$\mathcal{D}_{a}^{\prime}$ =$=\left\{(\bm{x}_{i}+\bm{r}_{i},\bm{y}_{i})|(\bm{x}_{i},\bm{y_{i}})\in\mathcal{D}_{a},\bm{r}_{i}\in\mathcal{R}\right\}_{i=1}^{|\mathcal{D}_{a}|},$
where $(\bm{x}_{i},\bm{y}_{i}),\bm{r}_{i}$ are the $i$-th element in $\mathcal{D}_{a}$ and $\mathcal{R}$. Inspired by [56], the adversary can adopt adversarial attack techniques [27, 70] to achieve the forgery attack.
@@
and $\bm{y}$ is the ground-truth label of $\bm{x}$. To attack IF-based methods, the adversary can set $\bm{y}_{t}=\bm{y}$ to increase the prediction confidence. In contrast, the adversary can set $\bm{y}_{t}\neq\bm{y}$ to attack EF-based methods. Notably, slight differences exist among the implementations of forged attacks against different dataset auditing methods, but they all follow the same strategy. We include the detailed descriptions in Appendix -D.
Arguably, the adversary can apply any targeted adversarial attacks to achieve such a goal. Specifically, we implement 5 different forgery attacks, as shown in Table I.
TABLE I: A list of attacks against dataset auditing implemented in DATABench. The ‘Data’ column shows whether they need auxiliary data for attacks.
@@
Gaussian Filter, Median Filter, Wavelet Filter Evasion Removal Pre-processing
@@
Adversarial Training Evasion Removal Training
Adaptively Splitting Poisoned Dataset (ASD) [26] Evasion Removal Training
@@
KNN-based Outlier Detection, SVM-based Outlier Detection Evasion Detection Post-training
FGSM [27], PGD [58], UAP [63] TIFGSM [17], VNIFGSM [82] Forgery / /
TABLE II: Results of evasion attacks in the preprocessing phase. We highlight cells corresponding to failed audits in orange and those indicating successful audits in blue. (Better viewed in color)
@@
V DATABench: Benchmarking Attacks against Dataset Auditing
V-A Overview of DATABench
In this section, we present the overview of our dataset auditing attack benchmark DATABench. Notably, DATABench aims to cover both simple and advanced attack methods. Table I shows the implemented attacks in DATABench. We implement a total of 17 evasion attacks, including 6 attacks in the pre-processing phase, 4 attacks in the training phase, and 7 attacks in the post-training phase. Naive attacks such as pruning and quantization are not included since a wide range of papers have demonstrated they are ineffective [49, 18, 41, 53]. Besides, we implement 5 different forgery attacks. Detailed descriptions of these attacks are in Appendix -B.
In DATABench, we implement 9 representative dataset auditing methods, including 4 IF-based methods and 5 EF-based methods. For IF-based methods, besides Dataset Inference (DI) [60] and Data-use Auditing (DUA) [41], which are designed for dataset auditing, we also adopt two representative MIAs as auditing methods. We select a classic MIA proposed by Shokri et al. [77] (denoted as MIA) and the state-of-the-art (SOTA) MIA named Rapid [35]. For EF-based methods, we evaluate DVBW [53], UBW (including UBW-P and UBW-C) [49], ZeroMark [29], and Domain Watermark (DW) [31]. The detailed settings of these dataset auditing methods can be found in Appendix -A.
TABLE III: Results of evasion attacks in the training phase. We highlight cells corresponding to failed audits in orange and those indicating successful audits in blue. (Better viewed in color)
@@ V-A Overview of DATABench
TABLE IV: Results of evasion attacks in the post-training phase. We highlight cells corresponding to failed audits in orange and those indicating successful audits in blue. (Better viewed in color)
@@
V-B Experimental Settings
Models and Datasets. In the main experiments, we adopt the widely-used CIFAR-10 [46] and ResNet-18 [32]. The reason for using this model and dataset is that it is a simple scenario for dataset auditing, while a challenging one for attacks [18]. Experiments on other models and datasets such as Vision Transformer [61] and ImageNet [16] can be found in Appendix -E. We further split the training data: 70% is the public release dataset (protected), 15% is the defender’s unpublished audit auxiliary dataset, and 15% is the adversary’s unpublished attack auxiliary dataset. Both auxiliary sets are assumed to be trusted and clean (unwatermarked). We train the model from scratch on the victim dataset. The experiments in the fine-tuning scenario are in Appendix -F.
Attack Settings. In our experiments, we evaluate existing dataset auditing methods on all the attacks in DATABench. Given the significant impact of data augmentation and regularization on model utility, we adopt them as default settings. Besides, we consider two variants for adversarial training attacks. ‘AdvTraining-FGSM’ simply uses FGSM [27] to generate the adversarial examples for training, and ‘AdvTraining-Hybrid’ composes three different adversarial attacks [27, 58, 7]. We also test two different privacy budgets ($\varepsilon=32$ or $64$) in DP-SGD to simulate different levels of privacy protection. Detailed descriptions and settings can be found in Appendix -B.
Dataset Auditing Settings. In our experiments, we consider the ideal situation for dataset auditing. If a dataset auditing method needs to train auxiliary models, we assume that the data owner uses the same model architecture as the adversary. It ensures that we evaluate the worst scenario for the adversary. Detailed settings can be found in Appendix -A.
@@ V-B Experimental Settings
  • Score: Score is the average confidence value outputted by the attack models in MIA and Rapid. A higher score indicates that the data is more likely to be a member (trained instance) of the training dataset, and the score $>0.5$ indicates a signal of a ‘trained’ dataset.
@@
  • WSR: WSR (Watermark Success Rate) measures the percentage of validation data getting the expected prediction, which is defined differently depending on the method: a specific label (DVBW/ZeroMark), an incorrect label (UBW-P/C), or the correct label (DW). We mark WSR $<25\%$ as a signal of a ‘non-trained’ dataset.
@@
V-C Results of Evasion Attacks
Tables II-IV show the results of evasion attacks in the pre-processing, training, and post-training phases, respectively. We summarize our key findings as follows.
@@ V-C Results of Evasion Attacks
A key finding is that dataset auditing methods relying on the embedding of strong external features, exemplified by DVBW and UBW-P in our experiments, exhibit relatively higher robustness. Their robustness is particularly exhibited against challenging training-phase attacks like adversarial training and DP-SGD. The two methods are poison-label methods that necessitate changing the labels of part of the samples in the dataset. However, even these methods are not immune to evasion attacks, showing vulnerabilities to specific pre-processing data transformations ($e.g.$, WaveletFilter, AutoEncoder) and post-training inference manipulations ($e.g.$, Reprogramming). Besides, other EF-based methods tested using mild external features ($i.e.$, UBW-C, ZeroMark, DW) displayed considerably less robustness across various attack stages.
@@
TABLE V: Results of the hybrid evasion attacks. We highlight cells corresponding to failed audits in orange and those indicating successful audits in blue. (Better viewed in color)

### v2

@@ 3.3 Forgery Attacks
$\mathcal{D}_{a}^{\prime}$ $=\left\{(\bm{x}_{i}+\bm{r}_{i},\bm{y}_{i})|(\bm{x}_{i},\bm{y_{i}})\in\mathcal{D}_{a},\bm{r}_{i}\in\mathcal{R}\right\}_{i=1}^{|\mathcal{D}_{a}|},$
where $(\bm{x}_{i},\bm{y}_{i}),\bm{r}_{i}$ are the $i$-th element in $\mathcal{D}_{a}$ and $\mathcal{R}$. Inspired by [57], the adversary can adopt adversarial attack techniques [29] to achieve the forgery attack.
@@
and $\bm{y}$ is the ground-truth label of $\bm{x}$. To attack IF-based methods, the adversary can set $\bm{y}_{t}=\bm{y}$ to increase the prediction confidence. In contrast, the adversary can set $\bm{y}_{t}\neq\bm{y}$ to attack EF-based methods. Notably, slight differences exist among the implementations of forged attacks against different dataset auditing methods, but they all follow the same strategy. We include the detailed descriptions in Appendix D.
Arguably, the adversary can apply any targeted adversarial attacks to achieve such a goal. Specifically, we implement 5 different forgery attacks, as shown in Table 1.
Table 1: A list of attacks against dataset auditing implemented in DATABench. The ‘Data’ column shows whether they need auxiliary data for attacks.
@@
Gaussian Filter, Median Filter,
Wavelet Filter
Evasion Removal Pre-processing
@@
Adversarial Training [80] Evasion Removal Training
Adaptively Splitting Poisoned
Dataset (ASD) [28]
Evasion Removal Training
@@
KNN-based Outlier Detection,
SVM-based Outlier Detection
Evasion Detection Post-training
FGSM [29], PGD [59], UAP [64]
TIFGSM [19], VNIFGSM [82]
Forgery / /
Table 2: Results of evasion attacks in the preprocessing phase. We highlight cells corresponding to failed audits in orange and those indicating successful audits in blue. (Better viewed in color)
@@
4 DATABench: Benchmarking Attacks against Dataset Auditing
4.1 Overview of DATABench
In this section, we provide an overview of our dataset auditing attack benchmark, DATABench. This work primarily focuses on dataset auditing for image classification models, as most existing studies concentrate on this domain and cannot be readily extended to other modalities or tasks. A discussion on dataset auditing for generative models is deferred to Section 5.2. In particular, our DATABench is designed to encompass both basic and advanced attack strategies. Table 1 shows the implemented attacks in DATABench. We implement a total of 17 evasion attacks, including 6 attacks in the pre-processing phase, 4 attacks in the training phase, and 7 attacks in the post-training phase. We exclude naive strategies such as pruning and quantization, as prior studies have consistently shown their ineffectiveness [51, 20, 42, 54]. Besides, we implement 5 different forgery attacks. Detailed descriptions of these attacks are in Appendix B.
In DATABench, we implement 9 representative dataset auditing methods, including 4 IF-based methods and 5 EF-based methods. For IF-based methods, besides Dataset Inference (DI) [61] and Data-use Auditing (DUA) [42], which are designed for dataset auditing, we also adopt two representative MIAs as auditing methods. We select a classic MIA proposed by Shokri et al. [77] (denoted as MIA) and the state-of-the-art (SOTA) MIA named Rapid [36]. For EF-based methods, we evaluate DVBW [54], UBW (including UBW-P and UBW-C) [51], ZeroMark [31], and Domain Watermark (DW) [33]. The detailed settings can be found in Appendix A.
Table 3: Results of evasion attacks in the training phase. We highlight cells corresponding to failed audits in orange and those indicating successful audits in blue. (Better viewed in color)
@@ 4.1 Overview of DATABench
Table 4: Results of evasion attacks in the post-training phase. We highlight cells corresponding to failed audits in orange and those indicating successful audits in blue. (Better viewed in color)
@@
4.2 Experimental Settings
Models and Datasets. In the main experiments, we adopt the widely used CIFAR-10 [47] dataset and the ResNet-18 [34] architecture. This setting, although simple from an auditing perspective, presents substantial challenges for adversaries, making it well-suited for identifying vulnerabilities of existing methods. To further demonstrate the generalizability of our findings, we also conduct experiments on other models and datasets, including Vision Transformer [62] and ImageNet [17], in Appendix G. We further split the training data: 70% is the public release dataset (protected), 15% is the defender’s unpublished audit auxiliary dataset, and 15% is the adversary’s unpublished attack auxiliary dataset. Both auxiliary sets are assumed to be trusted and clean (unwatermarked). We train the model from scratch on the victim dataset. The experiments in the fine-tuning scenario are in Appendix H.
Attack Settings. In our experiments, we evaluate existing dataset auditing methods against all attacks included in DATABench. Since data augmentation and regularization substantially affect model utility, we adopt them as default settings. For adversarial training attacks, we consider two variants: AdvTraining-FGSM, which leverages FGSM [29] to generate adversarial examples, and AdvTraining-Hybrid, which combines three attack methods [29, 59, 8]. For DP-SGD, we examine two privacy budgets ($\varepsilon=32$ and $64$) to simulate different levels of privacy protection. Detailed descriptions and configurations are provided in Appendix B. To ensure both simplicity and fairness, we adopt the most classical or their default configurations for each attack. We observe that this might introduce model utility degradation in certain scenarios, which we further discuss in Appendix E.
Dataset Auditing Settings. In our experiments, we consider the ideal situation for dataset auditing. If a dataset auditing method needs to train auxiliary models, we assume that the data owner uses the same model architecture as the adversary. It ensures that we evaluate the worst scenario for the adversary. Detailed settings can be found in Appendix A.
@@ 4.2 Experimental Settings
  • Score: Score is the average confidence value outputted by the attack models in MIA and Rapid. A higher score indicates that the data is more likely to be a member (trained instance) of the training dataset, and the score >0.5>0.5 indicates a signal of a ‘trained’ dataset.
@@
  • WSR: WSR (Watermark Success Rate) measures the percentage of validation data getting the expected prediction, which is defined differently depending on the method: a specific label (DVBW/ZeroMark), an incorrect label (UBW-P/C), or the correct label (DW). We mark WSR $25\%$ as a signal of a ‘non-trained’ dataset.
@@
4.3 Results of Evasion Attacks
Tables 2-4 show the results of evasion attacks in the pre-processing, training, and post-training phases, respectively. We also explore advanced attacks by combining different attacks (details in Appendix F). Our key findings are as follows.
@@ 4.3 Results of Evasion Attacks
A key finding is that dataset auditing methods relying on the embedding of strong external features ($e.g.$, DVBW and UBW-P) in our experiments exhibit relatively higher robustness. It is particularly exhibited against challenging training-phase attacks like adversarial training and DP-SGD. The two methods are poison-label methods that necessitate changing the labels of part of the samples in the dataset. However, even these methods are not immune to evasion attacks, showing vulnerabilities to specific pre-processing data transformations ($e.g.$, WaveletFilter, AutoEncoder) and post-training inference manipulations ($e.g.$, Reprogramming). Besides, other EF-based methods tested using mild external features ($i.e.$, UBW-C, ZeroMark, DW) displayed considerably less robustness across various attack stages.
Moreover, IF-based methods generally exhibit moderate robustness, with DUA demonstrating relatively higher robustness. DUA enhances auditing effectiveness by injecting additional noise into the dataset. This mechanism improves robustness but also makes the audit results more susceptible to forgery. We will further discuss the details in Section 4.4.
@@
Table 5: Results of forgery attacks. We highlight cells corresponding to failed audits in orange and those indicating successful audits in blue. (Better viewed in color)
Type Auditing Method Metric No Attack FGSM PGD UAP TIFGSM VNIFGSM
IF-based MIA Score ($\downarrow$) 0.54 0.66 0.72 0.54 0.71 0.71
Rapid Score ($\downarrow$) 0.22 0.79 0.89 0.30 0.86 0.88
DI p-value ($\uparrow$) 0.01 $10^{-13}$ $10^{-5}$ $10^{-23}$ 0.97 0.01
Diff ($\downarrow$) 0.79 1.35 0.98 1.61 -0.81 0.09
DUA Cost (logits) (%, $\uparrow$) 100.00 0.13 0.13 0.18 0.13 0.24
Cost (label) (%, $\uparrow$) 100.00 0.13 0.13 0.24 0.13 0.24
EF-based DVBW p-value ($\uparrow$) 1.00 0.00 0.00 0.00 0.00 0.00
WSR (%, $\downarrow$) 10.01 29.92 98.37 65.60 69.77 96.77
UBW p-value ($\uparrow$) 1.00 0.00 0.00 0.00 0.00 0.00
WSR (%, $\downarrow$) 11.73 86.55 100.00 71.08 95.69 100.00
ZeroMark p-value ($\uparrow$) 1.00 1.00 1.00 0.99 0.01 1.00
WSR (%, $\downarrow$) 10.01 29.93 98.44 68.28 71.68 96.88
DW WSR (%, $\downarrow$) 11.73 67.00 89.40 66.37 73.72 89.36
4.4 Results of Forgery Attacks
For forgery attacks, we assume that the adversary wants to make their own dataset ($i.e.$, the attack auxiliary dataset) to be falsely identified as the ‘trained dataset’ of other independent models. We use the model trained on the unmarked release dataset as the independent model. The results of white-box forgery attacks are in Table 5, and the experiments in the black-box scenario can be found in Appendix I.
Takeaway 4: Existing dataset auditing methods are overall vulnerable to forgery attacks.
Takeaway 5: Dataset auditing methods that involve more complex decision-making and auditing processes are more resistant to forgery.
4.5 Attack Overhead
4.6 Summary of Evaluated Methods
To provide a holistic comparison of the evaluated dataset auditing methods, we summarize their relative performance across five key dimensions: Robustness, Distinctiveness, Fidelity, Stealthiness, and Efficiency according to the design objectives in Section 2.4. Figure 4 represents this comparison using a radar chart, where scores closer to the outer edge (1.0) indicate better performance in that dimension.
5 Discussion
5.1 Related Works
Since dataset auditing in DL is an emerging topic, there are only a limited number of surveys or benchmarks that have explicitly explored this domain. Asswad et al.[2] highlighted the importance of protecting data ownership and reviewed representative related techniques. Chandrasekaran et al.[10] provided a comprehensive study on machine learning governance, including data ownership in machine learning. Recently, Du et al. [20] presented the first systematic SoK regarding dataset auditing, primarily focusing on surveying existing solutions. From the broader perspective of AI copyright protection, the majority of existing work focuses on watermarking techniques for models [69], generated content [70, 88], and traditional multimedia [55]. Lukas et al. [58] presented a toolbox to evaluate the robustness of DNN watermarking, which is fundamentally different from dataset auditing.
5.2 Dataset Auditing for Generative Models
Given these challenges, existing auditing methods for generative models are not maturely applicable and may yield unstable or misleading results when evaluated under adversarial settings. Therefore, we caution that rushing into adversarial robustness evaluations in such an ill-defined setting without reliable baselines may misrepresent the true difficulty of the problem or misguide the community toward premature conclusions. Nevertheless, arguably, our framework and attack methodology, structured by three processing stages ($i.e.$, pre-processing, training, post-training) and three attack strategies ($i.e.$, decoupling, removal, detection), are data-agnostic and could guide evaluations across different modalities and tasks. We leave this as an important direction for future work.
5.3 Potential Future Direction
Future Direction for Dataset Auditing. The demonstrated vulnerability of existing dataset auditing methods underscores an urgent need for fundamentally more robust and distinctive techniques. Future work may focus on developing novel auditing mechanisms resilient to these attacks. (1) For evasion, this could involve exploring IFs that are harder to decouple from the model utility. For EF-based methods, future works may seek EFs that are deeply integrated into the model’s core functionality rather than easily removable or detectable trigger patterns. A certified robust mechanism is also urgently needed. (2) For forgery, our empirical study suggests that complicating the auditing process may lead to better resistance. Consequently, exploring more complex methods, such as multi-bit watermarks [75] or cryptography [27], may be a promising future direction to improve dataset auditing.
Limitations and Future Directions for Evaluating Dataset Auditing. While DATABench provides a comprehensive benchmark, there are still some limitations. First, this paper primarily focuses on dataset auditing regarding image datasets since they are currently the most widely studied and feasible methods. There have been some preliminary works studying auditing datasets of other modalities, such as text [42, 84], point clouds [83], and audio [52]. Providing a comprehensive evaluation of these works may be an important future work. Second, this paper does not cover all possible scenarios of dataset usage in DL. We primarily evaluate the two most important and widely adopted settings: training from scratch (Section 4) and fine-tuning (Appendix H). Third, DATABench does not thoroughly explore finer-grained threat models, such as adversaries with varying levels of capabilities (an adversary with very limited computational resources or data). An exhaustive study of the attacker’s capabilities would be a meaningful direction for future work.
6 Conclusion
This paper presented a comprehensive adversarial evaluation of dataset auditing in Deep Learning (DL) systems, addressing the critical gap between existing dataset auditing methods and their resilience against dedicated attacks. We introduced a novel internal feature and external feature taxonomy to classify auditing techniques, formally defined evasion and forgery attack objectives, and proposed systematic attack strategies to realize them. These contributions led to the development of DATABench, a benchmark featuring 17 evasion and 5 forgery attacks. DATABench also provides an extensible toolbox to design new dataset auditing methods and attacks. This toolbox can also facilitate the unified evaluation of auditing methods. Our empirical experiments of 9 representative auditing methods against DATABench revealed some critical findings, $e.g.$, none of them demonstrated sufficient robustness or distinctiveness under adversarial settings. These results highlight the significant vulnerability of current dataset auditing methods and underscore the urgent need for developing novel and adversarially resilient dataset auditing methods.
Ethical Considerations
Appendix A Detailed Introduction to Dataset Auditing implemented in DATABench
Shokri et al. (MIA) [77]. Shokri et al. proposed the first membership inference attack (MIA), leveraging models’ differential behavior ($e.g.$, higher confidence) on members versus non-members. The method trains multiple ‘shadow models’ to generate labeled prediction vectors (member/non-member), which then train a final classifier to distinguish target model members. We used 10 shadow models in our experiments.
Rapid [36]. Rapid is a SOTA MIA addressing inherent errors in difficulty calibration methods of reference-based attacks. It combines the original membership signal ($e.g.$, loss) with the calibrated score from reference models. Rapid trains a scoring model mapping both scores to a final membership prediction, learning to correct limitations of relying solely on calibration. We used 4 reference models in our experiments.
Dataset Inference (DI) [61]. DI assesses if a suspicious model shows higher prediction margins for the owner’s data points. It generates feature embeddings capturing this margin from the suspicious model using white-box or black-box methods. A confidence regressor, pre-trained on the owner’s model and data, scores these embeddings. Finally, a statistical hypothesis test aggregates scores to determine if the suspicious model used the owner’s dataset.
Data-use Auditing (DUA) [42]. DUA is a general dataset auditing framework. The owner creates two marked versions of each datum (preserving utility, maximizing difference), publishing one and hiding the other. Then, it uses a black-box Membership Inference (MI) score function and compares scores on the published versus hidden versions (contrastive MI). A sequential hypothesis test then determines if the published data was used for training.
DVBW [54]. DVBW proposes the first and a general framework for backdoor-based dataset auditing. Specifically, DVBW utilizes poison-label backdoor attacks for dataset auditing and presents a hypothesis-test-based verification method. In our experiments, the target label is set to 1.
UBW-P/C [51]. UBW is proposed for harmless and stealthy dataset copyright protection using untargeted backdoors. UBW has two variants, UBW-P and UBW-C. UBW-P utilizes an untargeted poison-label backdoor attack that changes the labels of the poisoned samples randomly. In contrast, UBW-C adopts bi-level optimization to craft clean-label ($i.e.$, the labels of the watermarked samples do not change) samples. Finally, UBW also utilizes a hypothesis test to complete dataset auditing.
ZeroMark [31]. ZeroMark is proposed for secure dataset auditing without disclosing watermarks. Inspired by the intrinsic property that boundary gradients of watermarked DNNs align with the watermark pattern, ZeroMark generates boundary versions of benign samples and calculates their boundary gradients under a label-only black-box setting. It then performs a hypothesis test based on the cosine similarity between the gradients and the secret watermark pattern to achieve auditing.
Domain Watermark (DW) [33]. DW is proposed for harmless dataset auditing by identifying a hardly-generalized domain. Instead of causing misclassification, DW ensures watermarked models correctly classify specific ‘hard’ samples misclassified by benign models. DW employs bi-level optimization to generate the domain and craft visually indistinguishable clean-label modified samples for stealth.
Appendix B Details of Evasion Attacks
Data Augmentation. Data augmentation is a widely-used approach to apply transformations to data to increase its diversity. In our experiments, we apply random cropping, resizing, and flipping as the augmentation methods.
Data Synthesis. Data synthesis involves generating artificial data, typically using generative models trained on the original dataset, to train the suspect model, thereby obscuring the direct link to the protected data. In our experiments, we use the DDIM diffusion model [37] as the generative model.
Denoising Filters. Denoising filters apply traditional image processing filters ($e.g.$, Gaussian, Median, Wavelet) to potentially remove noise-like external features (EF) embedded for auditing, effectively cleaning the data before model training. In our experiments, we set the kernel size to be 3.
Autoencoder-based Denoising. Autoencoder-based Denoising employs a trained autoencoder model to reconstruct the data and remove EFs. In our experiments, we adopt a 6-layer CNN-based autoencoder model.
Regularization. Regularization refers to the weight decay during training to prevent the model from overfitting. In our experiments, we use the classic $\ell_{2}$ regularization.
DP-SGD. DP-SGD adds calibrated noise during training to obscure the model’s dependence on specific training examples. In our experiments, we apply the Gaussian mechanism and test two different privacy budgets, 32 and 64.
Adversarial Training. Adversarial Training augments the training data with adversarially perturbed examples. We use two different implementations of adversarial training. AdvTraining-FGSM uses solely the FGSM method [29] to generate adversarial examples while AdvTraining-Hybrid adopts three methods, FGSM [29], PGD [59], and C&W [8].
Adaptive Splitting Poisoned Dataset (ASD) [28]. ASD is a training-time backdoor defense. ASD employs loss-guided splitting and meta-learning-inspired splitting to dynamically update two separate data pools. By leveraging the clean and poisoned data pools obtained through this adaptive partitioning, ASD effectively mitigates backdoors during training.
Noisy Output&Feature. Noise is injected into the model’s intermediate feature activations or final output predictions during inference to mask the subtle signals. In our experiments, we use Gaussian noise and the standard deviation is set to 0.5.
Randomized Smoothing (RS). RS smooths the model’s predictions by averaging outputs over multiple noisy input perturbations. We set the number of perturbed samples to 30.
Reprogramming [13]. Reprogramming modifies the model’s inference process, often using learnable input and output transformations, to neutralize the specific abnormal behaviors. In our experiments, we use a 10-layer U-Net as the input transformation module and train the U-Net for 10 epochs.
SCALE-UP [32]. SCALE-UP is an effective black-box, input-level backdoor detection method. SCALE-UP identifies and filters malicious test samples by analyzing their prediction consistency during the pixel-wise amplification process.
Outlier Detection (OD). OD identifies input samples that are statistically different from typical data (potentially EF triggers), allowing the adversary to return a non-standard or random output to evade detection. In DATABench, we use two different models, $i.e.$, KNN and SVM, to achieve OD.
Appendix C Details of Forgery Attacks
In forgery attacks, we implement different targeted adversarial attacks to craft fake ‘trained’ samples. Specifically, for DUA, DVBW, ZeroMark, and DW. We set the target label to 4. For UBW-P/C, since UBW tests whether the testing samples are misclassified (without specifying the target label), we implement the forgery attacks using untargeted methods. For DI, MIA, and Rapid, our objective is to increase the output confidence of these forged samples. Therefore, we apply a reverse untargeted attack ($i.e.$, in contrast to the optimization direction of untargeted attacks) to maximize confidence.
Appendix D Other Detailed Experimental Settings
As default settings, we train ResNet-18 on CIFAR-10 from scratch for 90 epochs with a learning rate of 0.03. The batch size is set to 128. Moreover, since batch normalization in ResNet-18 does not support DP-SGD, we alter these layers with group normalization. For EF-based methods, we use the Blended method [12], which applies a weighed blend of a specific trigger pattern and the original image, with the trigger pattern assigned a weight of 0.2. The experiments are conducted using at most 4 NVIDIA RTX 3090 GPUs.
Appendix E Discussion of the Model Utility Degradation
To ensure a fair, simple, and reproducible evaluation, we adopt the classical or standard/default configurations for each attack implemented in our DATABench. While this choice may lead to some degradation in model utility under certain scenarios, we argue that the key findings of this paper remain valid for the following two main reasons.
  • Some evasion techniques are inherently designed with objectives that conflict with maximizing model utility. A representative example is DP-SGD, which deliberately injects noise into the training process to provide certified privacy guarantees. As demonstrated in our experiments, this mechanism substantially reduces test accuracy while effectively evading auditing methods. In such cases, the degradation in utility is an intentional and acceptable trade-off for achieving the desired level of privacy.
  • Arguably, a moderate reduction in a model’s final utility does not diminish the value gained from the unauthorized use of a dataset. Even if an adversary’s model attains lower accuracy after applying an evasion attack, the protected dataset has still been exploited and has contributed to the model’s training and development. Thus, the core concern of unauthorized data usage—whether in terms of privacy breaches or copyright infringements—remains unresolved.
Table 6: Results of the hybrid evasion attacks. We highlight cells corresponding to failed audits in orange and those indicating successful audits in blue. (Better viewed in color)